OpenStudy (kittiwitti1):

Limiting reagent/factor problem. We have 87 A, 83 B, 168 C and 328 D. The requirements are to create a pack of 1 A, 1 B, 2 C, and 4 D. What is the limiting reagent?

OpenStudy (openstudygirl2):

ok... ummmmm

OpenStudy (openstudygirl2):

ok try first work it out yourself by trying what u think is right ok try that first

OpenStudy (openstudygirl2):

... tell me when ur done

OpenStudy (openstudygirl2):

@kittiwitti1

OpenStudy (openstudygirl2):

uhhhh ok then

OpenStudy (openstudygirl2):

msg me if your back and i will help you

OpenStudy (kittiwitti1):

I can't work it out myself :\

OpenStudy (openstudygirl2):

ok then take a random guess

OpenStudy (kittiwitti1):

I know how to calculate with 2-3 variables but not 4

OpenStudy (openstudygirl2):

ok then if u were soooo good at math but this one question you could not answer and then a man told you that the question is simple you just have to take a random pick and that will be the answer then what would you say the answer is

OpenStudy (openstudygirl2):

ok @kittiwitti1

OpenStudy (openstudygirl2):

ok then if u were soooo good at math but this one question which was this question you could not answer and then a man told you that the question is simple you just have to take a random pick and that will be the answer then what would you say the answer is

OpenStudy (kittiwitti1):

Uh... it doesn't work like that... also, your question is getting help

OpenStudy (openstudygirl2):

oh ok

OpenStudy (openstudygirl2):

ummm my dad is home and its my next sub so i rly need to do my next sub sorry um but keep asking people and then fan and medal them byeee

OpenStudy (kittiwitti1):

yeah I understand, go do your work.

OpenStudy (kittiwitti1):

@zzr0ck3r ?

OpenStudy (zzr0ck3r):

i.e. he can't help you because he does not know how.... and randomly guessing is a HORRIBLE idea. I unfortunately have no idea what you are talking about in this question.

OpenStudy (kittiwitti1):

I believe limiting reagent refers to which of the variables limits the amount of product produced by the other variables.

OpenStudy (kittiwitti1):

If that helps...

OpenStudy (mathmate):

Question was: `Limiting reagent/factor problem. We have 87 A, 83 B, 168 C and 328 D. The requirements are to create a pack of 1 A, 1 B, 2 C, and 4 D. What is the limiting reagent?` @kittiwitti1 said `I believe limiting reagent refers to which of the variables limits the amount of product produced by the other variables.` For each agent/material, divide each available value by the required value to find the factor of that agent. Use the minimum value of the factors and round it down as the factor of all agents. Suppose the available agents are 23A,27B,39C,80D, then 23A/1A=23 27B/1B=27 39C/2C=19.5 80D/4D=20 The minimum value is 19.5, the greatest integer not exceeding 19.5 is 19. So 19 can be used as the factor, and C is therefore the limiting agent, because everything else has excess.

OpenStudy (kittiwitti1):

That was very lengthy but I understood. Thank you for taking the time @mathmate :)

OpenStudy (mathmate):

Agreed. I prefer to be a little more detailed to help you understand. In math and science, understanding and applying what you learned is the secret to learning future topics.